# Machine learning for hydrodynamic stability

### v2

@@ Abstract.
A machine-learning strategy for investigating the stability of fluid flow problems is proposed herein. The objective is to exploit shallow neural networks for a binary classification task to distinguish between bifurcated and non-bifurcated states. The goal is to provide a simple yet robust methodology to find a nonlinear mapping from the parametric space to a bifurcating indicator identifying the region of interest. The computational procedure is demonstrably robust and does not require parameter tuning. The essential feature of the strategy is that the computational solution of the Navier–Stokes equations is a reliable proxy for laboratory experiments investigating sensitivity to flow parameters. The applicability of our bifurcation detection strategy is demonstrated by an investigation of two classical examples of flow instability associated with thermal convection. The codes used to generate and process the labelled data are available on GitHub.
Acknowledgements. This work was supported by EPSRC grants EP/W033801/1 and EP/Y028783/1.
@@ 1. Introduction
The issue of hydrodynamic stability has been extensively studied over the last century. Excellent books on the topic include Joseph [14], Landau & Lifshitz [15], Drazin & Reid [6] and Schmid & Henningson [18]. The graduate textbook by Drazin [5] provides a perfect introduction to the subject. Linearizations for small perturbations of a base (steady) flow are the starting point for developing mathematical theory and often provide results that match with laboratory experiments. Classical linear stability analysis assumes exponential behaviour of infinitesimal perturbations and leads to characteristic eigenvalue problems that can be analysed theoretically (following in the footsteps of Lord Rayleigh). They can also be explored computationally by discretising the governing partial differential equations (PDEs) linearised about the base flow and solving a high-dimensional eigenvalue problem to track the critical eigenvalues. See Cliffe et al. [4] for a comprehensive review and Wubs & Dijkstra [23, Chap. 6] for a contemporary treatment.
The concept of “learning from data” has become mainstream over the past decade with activity in all areas of science. As noted by Strang [21] the central goal of machine learning is to construct a function that classifies the training data correctly, so it can generalise to unseen test data. Assessing the effectiveness of a rudimentary (shallow) neural network in correctly classifying hydrodynamic stability is our primary focus here.
@@
Starting from the pioneering paper of Raissi et al. [17], problems from fluid dynamics have featured prominently in papers exploring machine learning of solutions of partial differential equations. Efforts to train physics-informed neural networks (PINNs) in the context of fluid flow modelling have met with mixed success to date however. Wang et al. [22] identify the Stokes and Navier–Stokes equations as challenging problems that highlight some of the most prominent difficulties in training PINNs. Recent work of Pichi et al. [16] builds on earlier work by Chen et al. [3] in constructing a machine-learning strategy for learning the steady-state bifurcation structure of flow problems. The developed strategy is limited by a computationally expensive offline stage whereby a reduced basis set is computed from a singular-value decomposition of a (high-dimensional) snapshot matrix. The associated error in this dimensional reduction is difficult to quantify and limits the achievable accuracy when modelling fluid flow (because the decay of the singular values is so slow). In contrast, the methodology developed herein is based on direct simulation of the flow problem, where the effect of discretisation error can be monitored to ensure that it does not influence the resulting flow classification. Neural network training is compared with Gaussian process regression in a recent paper by Sousedik et al. [20] assessing the linear stability of two benchmark flow problems. The issues associated with the computation and careful tracking of eigenvalues are eschewed in our bifurcation detection strategy.
@@
Regarding the discretisation strategy, a simple method-of-lines approach that combines a stable finite element spatial approximation together with a similarly accurate implicit time discretisation is adopted herein. One feature distinguishing our strategy from that of others is the use of self-adaptive time stepping that is designed to “follow the physics”. The benefit of this is that instability of the computed flow does not need to be explicitly excited. In singular cases (multiple steady-state flows) the onset of instability is triggered by computer round-off much as it would be triggered in physical experiments by geometrical imperfections in the laboratory apparatus. The upshot however, is that stable linear solver strategies are essential if we are to respect the singularity or ill-conditioning of the linearised algebraic equation systems. These issues are explored in the context of flow problem with a symmetry-breaking pitchfork flow bifurcation in section 3. The applicability of our bifurcation detection strategy is demonstrated in section 4, where we present results for two examples of flow instability associated with thermal convection that have been partially studied previously.
@@ 2. Bifurcation classification using a neural network
The two distinct flow problems considered in this work are defined on a bounded spatial domain $\Omega\in\mathbb{R}^{2}$ and a temporal domain $\tau=(0,\infty)$. Given a parabolic PDE operator $L$ (typically representing the Navier–Stokes equations) and a vector $\lambda\in\mathbb{R}^{n}$ of parameters, the goal is to explore the bifurcation structure of flow solutions $u({\vec{x}},t;\lambda)$ satisfying
$L(u({\vec{x}},t;\lambda))&=0&&\text{in $\Omega\times\tau$},\\ u(\cdot,t;\lambda)&=g(t;\lambda)&&\text{on $\partial\Omega\times\tau$},\end{aligned}$ (1)
starting from a quiescent flow profile at time $t=0$.
The characterisation of the bifurcation structure of a given flow problem is a two-stage procedure. In the first stage, approximations $u_{h}$ of solutions $u$ will be computed by discretising (1) in space (using mixed finite element approximation) and in time (using adaptive time stepping driven by local error control), by sampling parameter vectors $\lambda$ in the neighbourhood of the bifurcation boundary. [^1] This step requires a priori knowledge of the bifurcation phenomenon. While this can often be guessed by physical considerations or from laboratory experiments, determining suitable sample points would pose a challenge in cases where the system behaviour is totally unknown. A classification test is then applied to the sampled data to determine whether the computed flow is representative of the base flow or else has undergone a bifurcation. The result of this is a set of labelled data
@@
where the label $\ell^{j}$ is either 0 or 1. A distinguishing feature of our time stepping strategy is that a bifurcated solution $(u_{h}^{j},1)$ typically evolves from the associated base state (this depends on the underlying bifurcation structure). Small perturbations of the flow solution are not explicitly introduced in the time stepping process. This supports our assertion that the computer can be viewed as a proxy for a corresponding laboratory experiment investigating sensitivity to flow parameters.
The second stage of our procedure is the generation of a surrogate model of the bifurcation structure. We will show that this can be done effectively by training a shallow neural network to classify the data. The machine-learning ingredients will be elementary: a shallow network will prove to be perfectly adequate and we do not need to carefully tune any hyperparameters. A random division of the data in $SS\SS$ into a training dataset and a test dataset will demonstrate that the machine-learned models of bifurcation structure are statistically valid. A representative surrogate generated with fixed PDE discretisation parameters can also be validated a posteriori, by sampling parameters close to the bifurcation boundary and comparing the model predictions with the labelled data that is generated for the same parameter values using a refined finite element mesh or a smaller time accuracy tolerance.
The neural network architecture was fixed throughout this study. The input to the network is the parameter vector $\lambda$ (two-dimensional in all cases discussed later) and the final output is a prediction of the corresponding label $\ell$. The network has a single hidden layer consisting of 32 neurons. Changing the width to 16 or to 64 neurons is not going to significantly affect the results. The activation function is the standard sigmoid. The output from the hidden layer is a two-dimensional vector $[z_{1},z_{2}]$ that is mapped onto a vector of probabilities $p=[p_{1},p_{2}]$ using a numerically stable softmax function
@@
In all cases discussed in later sections, the training of the network proved to be robust with respect to the initialisation of the weights and biases and did not require bespoke parameter tuning. [^2] The raw parameter data does need to be normalised however. This was done by subtracting the mean and dividing by the standard deviation of the individual components. A typical training run with ${m}=500$ data points with a fixed learning rate $\alpha=0.15$ is shown in Figure 1. The optimisation in this instance was 5000 passes through the dataset, taking a gradient descent step after every data point with gradients computed analytically by back-propagation. Looking more closely at the convergence history, we see that the output of the loss functional decreases non-monotonically initially with a reduction of an order of magnitude in the first 12 steps followed by a monotonic reduction by an additional factor of 2 (50%) after 36 steps. The rate of decrease of the loss functional is relatively slow thereafter: the relative reduction achieved in the final 4000 steps is close to 25%. The $\ell_{2}$ norm of the gradient vector when the training run ended was $1.49\times 10^{-7}$. The time taken for the complete training run was less than 4 minutes when run using the open-source Python code on a 2018 MacBook Pro.
Figure 1. Network training history for 500 data points (symmetric channel flow bifurcation).
The use of a neural network surrogate to model the stability of flow configurations with multiple parameters is completely novel as far as we are aware. The methodology will be shown to offer new insights into the assessment of the stability of the flow problems that feature in this study.
@@ 3. Validation of the classification strategy
The effectiveness of our classification strategy will be demonstrated in the first instance by studying a classical symmetry-breaking channel flow test problem and comparing results with those in the literature. Specifically, we compute approximations to the velocity $\vec{u}=[u_{x},u_{y}]$ and pressure $p$ satisfying the Navier–Stokes equations in a rectangular duct with a symmetric expansion. Details of the domain and a typical steady-state solution (below the critical Reynolds number with an unperturbed inflow) are shown in Figure 2. The associated steady-state boundary conditions are
@@
To be consistent with previous work, we will characterise the flow Reynolds number $\mathcal{R}=LU/\nu$ in terms of the kinematic viscosity $\nu$, the maximum speed at the inflow $U$ and the height of the channel at the outflow $L$. The symmetry of the base solution can be confirmed by observing that the two eddies downstream of the step in Figure 2 have the same reattachment length.
Refer to caption11222216161616
@@
[…]
@@
The specific flow problem is referred to as a 1:2 expansion in the computational study reported by Drikakis [7] and it exhibits a pitchfork bifurcation if the Reynolds number is increased past a critical value $\mathcal{R}^{*}$ of approximately 220. At the critical value the rightmost eigenvalue of the linearised Jacobian matrix (see Elman & Silvester [10]) changes from a real negative number (indicating linear stability) to a real positive number (instability). Just above the critical Reynolds number there are three steady solutions: the symmetric flow solution is unstable and there are two stable asymmetric solutions with the two recirculating eddies having a different length. One such solution computed at a Reynolds number $\mathcal{R}>\mathcal{R}^{*}$ is shown in Figure 3. The second steady state is the mirror image: the top and bottom eddies in Figure 3 are inverted.
@@
$\omega(t)=\int_{\Omega}\nabla\times\vec{u}(\cdot,t)\,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rm{d}\Omega}=\int_{\partial\Omega_{N}}\!\!\!\!u_{y}(\cdot,t)\,\rm{d}s,$ (7)
@@
An issue that it is important to appreciate is that classical eigenvalue analysis provides no information about the size of asymmetric perturbation needed to the initiate a transition to an asymmetric flow profile from an unstable symmetric flow pattern. Herein, to replicate a practical laboratory experiment, the influence of perturbation magnitude will be studied. Specifically, we aim to assess the symmetry or otherwise of an evolving flow generated with a distorted steady-state parabolic inlet profile,
@@
We generate data for two different sets of parameters $\lambda=(\mathcal{R},\delta)$. The first range of results is generated with 35 uniformly spaced Reynolds numbers between 209.6 and 225.7 and 10 logarithmically spaced perturbations between $10^{-16}$ and $10^{-3}$. The second range is generated with 30 uniformly spaced Reynolds numbers between 168.1 and 222.2 and 5 logarithmically spaced perturbations between $10^{-5}$ and $5\times 10^{-2}$. Combining these two data sets into a single set $SS\SS$, 281 data points were labelled as symmetric flows and 219 were classed as asymmetric.
@@
Figure 7. Symmetric step channel flow: data classification (left); surrogate predictions from trained neural network (right). Red dots correspond to $\ell^{j}$=0 and blue dots correspond to $\ell^{j}$=1.
To test the predictive capability of the surrogate model, the data set $SS\SS$ is randomly divided into a training set with 375 data points and a test set with 125 data points. The classification of the test data in one such division is shown in Figure 6. In this representative data split the surrogate model generated by the training data correctly classifies 123 of the test data points with 2 asymmetric flows incorrectly classified as being symmetric. The test was repeated multiple times. The average precision after 5 independent tests was 0.964, and the average F1 score was 0.969. The visualisation of the bifurcation boundary in Figure 7 is generated by training the neural network on 375 data points, sampling the parameter set at 50,000 points normally distributed around $\mathcal{R}=218$ and plotting the predicted classification at these points.
@@
Figure 8. Symmetric step channel flow: refined grid classification (left) and prediction accuracy (right). Red dots on the right are correct predictions, blue dots are incorrect.

### v4

@@ Abstract.
A machine-learning strategy for investigating the stability of fluid flow problems is proposed herein. The goal is to provide a simple yet robust methodology to find a nonlinear mapping from the parametric space to an indicator representing the probability of observing a bifurcated solution. The computational procedure is demonstrably robust and does not require parameter tuning. The essential feature of the strategy is that the computational solution of the Navier–Stokes equations is a reliable proxy for laboratory experiments investigating sensitivity to flow parameters. The applicability of our probabilistic bifurcation detection strategy is demonstrated by an investigation of two classical examples of flow instability associated with thermal convection. The codes used to generate and process the labelled data are available on GitHub.
Acknowledgements. This work was supported by EPSRC grants EP/W033801/1 and EP/Y028783/1. The authors would like to acknowledge the support provided by Research IT and the use of the Computational Shared Facility at The University of Manchester.
@@ 1. Introduction
The issue of hydrodynamic stability has been extensively studied over the last century. Excellent books on the topic include [Joseph(2014)], [Landau and Lifshitz(1987)], [Drazin and Reid(1981)] and [Schmid and Henningson(2001)]. Linearizations for small perturbations of a base (steady) flow are the starting point for developing mathematical theory and often provide results that match with laboratory experiments. Classical linear stability analysis assumes exponential behaviour of infinitesimal perturbations and leads to characteristic eigenvalue problems that can be analysed theoretically (following in the footsteps of Lord Rayleigh). They can also be explored computationally by discretising the governing partial differential equations (PDEs) linearised about the base flow and solving a high-dimensional eigenvalue problem to track the critical eigenvalues. See [Cliffe et al.(2000)] for a comprehensive review and [Wubs and Dijkstra(2023), Chap. 6] for a contemporary treatment. A practical issue is that classical eigenvalue analysis provides no information about the size of asymmetric perturbation needed to initiate a transition to an asymmetric flow profile from an unstable symmetric flow pattern. In this work, the influence of the perturbation magnitude will be assessed. The motivation for this is to provide insight into the relative difficulty of assessing flow stability by laboratory experiment. The concept of “learning from data” has become mainstream over the past decade with activity in all areas of science. Quoting from [Strang(2019)], the central goal of machine learning is to construct a function that classifies the training data correctly, so it can generalise to unseen test data. Assessing the effectiveness of a rudimentary neural network in correctly classifying hydrodynamic stability will be the focus of this study.
There is a bugeoning literature on machine learning techniques for bifurcation problems. State-of-the-art machine learning techniques, such as those discussed by [Szep et al.(2021)] or [Mogharabin and Ghadami(2025)] can efficiently determine multiple bifurcations in cases where the underlying dynamical system is low-dimensional. Julia software tools such as ModelingToolkit.jl can autodifferentiate ODE systems and when interfaced with BifurcationKit.jl, can map from right-hand sides to continuation and diagram construction. However, as noted in [Szep et al.(2021)], these sohisticated detection strategies may not scale effectively for partial differential equations where a large number of states arises from discretisation of the spatial variables. Other specialised machine learning methods such as those discussed in [Celletti et al.(2025)] can be used to classify different types of motions (regular or chaotic) from a time series or even be used to approximate Lyapunov exponents using neural networks [Mayora-Cebollero et al.(2025)]. Both of these tasks are significantly harder than the binary classification distinguishing between bifurcated and non-bifurcated states that is the focus of this study.
@@
Starting from the pioneering paper [Raissi et al.(2020)], problems from fluid dynamics have featured prominently in papers exploring machine learning of solutions of partial differential equations. Efforts to train physics-informed neural networks (PINNs) in the context of fluid flow modelling have met with mixed success to date however. [Wang et al.(2023)] identify the Stokes and Navier–Stokes equations as challenging problems that highlight some of the most prominent difficulties in training PINNs. Recent work [Pichi et al.(2023)] builds on earlier work [Chen et al.(2021)] in constructing a machine-learning strategy for learning the steady-state bifurcation structure of flow problems. The developed strategy is limited by a computationally expensive offline stage whereby a reduced basis set is computed from a singular-value decomposition of a (high-dimensional) snapshot matrix. The associated error in this dimensional reduction is difficult to quantify and limits the achievable accuracy when modelling fluid flow (because the decay of the singular values is so slow). In contrast, the methodology developed herein is based on direct simulation of the flow problem, where the effect of discretisation error can be monitored to ensure that it does not influence the resulting flow classification. Neural network training is compared with Gaussian process regression in a recent paper [Sousedík et al.(2022)] assessing the linear stability of two benchmark flow problems. The issues associated with the computation and careful tracking of eigenvalues are eschewed in our bifurcation detection strategy.
@@
Regarding the discretisation strategy, a standard method-of-lines approach that combines a stable finite element spatial approximation together with a similarly accurate implicit time discretisation is adopted herein. One feature distinguishing our strategy from that of others is the use of self-adaptive time stepping that is designed to “follow the physics”. The benefit of this is that instability of the computed flow does not need to be explicitly excited. In singular cases (multiple steady-state flows) the onset of instability is triggered by computer round-off much as it would be triggered in physical experiments by geometrical imperfections in the laboratory apparatus. The upshot however, is that stable linear solver strategies are essential if we are to respect the singularity or ill-conditioning of the linearised algebraic equation systems. These issues are explored in the context of flow problem with a symmetry-breaking pitchfork flow bifurcation in section 3. The applicability of our bifurcation detection strategy is demonstrated in section 4, where we present results for two examples of flow instability associated with thermal convection that have been partially studied previously.
The use of a neural network surrogate to model the stability of flow configurations generated by direct simulation of the Navier–Stokes equations is completely novel as far as we are aware. The methodology will be shown to offer new insights into the assessment of the stability of the specific flow problems that feature in this study.
@@ 2. Bifurcation classification using a neural network
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathscr{L}(u({\vec{x}},t;\lambda),p({\vec{x}},t;\lambda))}=0$ $\quad\text{in $\Omega\times\tau$},$ (1)
starting from a quiescent flow at time $t=0$, combined with physically realistic boundary conditions on $\partial\Omega\times\tau$.
The characterisation of the bifurcation structure of a given flow problem is a two-stage procedure. In the first stage, approximations $u_{h}$ of solutions $u$ will be computed by discretising (1) in space (using mixed finite element approximation) and in time (using adaptive time stepping driven by local error control), by sampling parameter vectors $\lambda^{j}$ in the neighbourhood of the bifurcation boundary. [^1] This step requires a priori knowledge of the bifurcation phenomenon. While this can often be guessed by physical considerations or from laboratory experiments, determining suitable sample points would pose a challenge in cases where the system behaviour is totally unknown. A classification test is then applied to the sampled data to determine whether the computed flow is representative of the base flow or else has undergone a bifurcation. The result of this is a set of labelled data
[…]
where the label $\ell^{j}$ is either 0 or 1. A distinguishing feature of our time stepping strategy is that a bifurcated flow solution $u_{h}^{j}$ typically evolves from the associated base state (this depends on the underlying bifurcation structure). Small perturbations of the flow solution are not explicitly introduced into the time stepping process. This supports our assertion that the computer can be viewed as a proxy for a corresponding laboratory experiment investigating sensitivity to flow parameters.
The second stage of our procedure is the generation of a surrogate model of the bifurcation structure. We will confirm that this can be done simply and effectively by training a shallow neural network to classify the data. The machine-learning ingredients will be elementary: a shallow network will prove to be perfectly adequate and we do not need to carefully tune any hyperparameters. A random division of the data in $S\SS$ into a training dataset and a test dataset will demonstrate that the machine-learned models of bifurcation structure are statistically valid. A representative surrogate generated with fixed PDE discretisation parameters can also be validated a posteriori, by sampling parameters close to the bifurcation boundary and comparing the model predictions with the labelled data that is generated for the same parameter values using a refined finite element mesh or a smaller time accuracy tolerance.
The neural network architecture was fixed throughout this study. The input to the network is the parameter vector $\lambda$ (two-dimensional in all cases discussed later) and the final output is a prediction of the corresponding label $\ell$. The network has a single hidden layer consisting of 32 neurons. Changing the width to 16 or to 64 neurons does not affect the results. The activation function is the standard sigmoid. The output from the hidden layer is a two-dimensional vector $z_{1},z_{2}]$ that is mapped onto a vector of probabilities $p=[p_{1},p_{2}]$ using a numerically stable softmax function
@@
Figure 1. Network training history for 105 data points (symmetric channel flow bifurcation).
In all cases discussed in later sections, the training of the network proved to be robust with respect to the initialisation of the weights and biases and did not require bespoke parameter tuning. [^2] The raw parameter data does need to be normalised however. This was done by subtracting the mean and dividing by the standard deviation of the individual components. A typical training run with m=105{m}=105 data points with a fixed learning rate $\alpha=0.15$ is shown in Figure 1. The optimisation algorithm consisted of 5000 passes through the entire dataset, taking a gradient descent step after every data point with gradients computed analytically by back-propagation. Looking closely at the convergence history, we can see that the output of the loss functional decreased monotonically with a reduction of an order of magnitude in the first 20 steps followed by a monotonic reduction by a further factor of 2 (50%) after 200 steps. The monotonic decrease of the loss functional is increasingly slow thereafter: the relative reduction achieved in the final 4000 steps was an additional factor of 10. The $\ell_{2}$ norm of the gradient vector when the training run ended was $3.16\times 10^{-35}$.
@@ 3. Validation of the classification strategy
The effectiveness of our classification strategy will be demonstrated in the first instance by studying a classical symmetry-breaking channel flow test problem and comparing results with classical stability results taken from the literature. Specifically, we compute approximations to the velocity $\vec{u}=[u_{x},u_{y}]$ and pressure $p$ satisfying the Navier–Stokes equations
${\partial{\vec{u}}\over\partial t}+{\vec{u}}\cdot\nabla{\vec{u}}-{1\over\mathcal{R}}\,\nabla^{2}{\vec{u}}+\nabla p$ $={\vec{0}},$ (6)
$\nabla\cdot{\vec{u}}$ $=0,$
in a rectangular duct with a symmetric expansion. Details of the domain and a typical steady-state solution (below the critical Reynolds number with an unperturbed inflow) are shown in Figure 2. To be consistent with previous work, we will characterise the flow Reynolds number $\mathcal{R}=LU/\nu$ in terms of the kinematic viscosity $\nu$, the maximum speed at the inflow $U$ and the height of the channel at the outflow $L$.
The associated steady-state boundary conditions are
@@
The symmetry of the base solution can be confirmed by observing that the two eddies downstream of the step in Figure 2 have the same reattachment length.
Refer to caption11221616
@@
The ifiss software package [Silvester et al.(2023)] is used to compute numerical solutions. The spatial discretization is effected using $Q_{2}$$P_{-1}$ (biquadratic velocity; discontinuous linear pressure) mixed approximation, see [Elman et al.(2014b), Section 3.3.1]. Labelled data is generated by discretising the flow domain into a uniform grid of 2112 square elements. [^3] The open source MATLAB software toolbox is described in [Elman et al.(2014a)]. The grid is defined by setting the discretization level to 5. The dimension $n_{x}$ of the resulting system of ordinary differential equations is 23,842. The (unstabilised) TR–AB2 integrator described in [Elman et al.(2014b), Sect. 10.2.3] is used for evolving the numerical solution. The time integration is started from quiescent flow $\vec{u}_{h}=0$ in the channel domain. The initial time step is 1e-9. To model the action of “turning on the tap” in a laboratory experiment, the inhomogeneous inflow boundary condition is smoothly ramped up. This is done by multiplying the steady-state inflow boundary profile by the lifting function $1-e^{-10t}$.
@@
The specific flow problem is referred to as a 1:2 expansion in the computational study described in [Drikakis(1997)]. The flow exhibits a pitchfork bifurcation if the Reynolds number is increased past a critical value $\mathcal{R}^{*}$ of approximately 220. At the critical value the rightmost eigenvalue of the linearised Jacobian matrix (see [Elman and Silvester(2018)]) changes from a real negative number (indicating linear stability) to a real positive number (instability). Just above the critical Reynolds number there are three steady solutions: the symmetric flow solution is unstable and there are two stable asymmetric solutions with the two recirculating eddies having a different length. One such solution computed at a Reynolds number $\mathcal{R}>\mathcal{R}^{*}$ is shown in Figure 3. The second steady state is the mirror image: the top and bottom eddies in Figure 3 are inverted.
@@
$\omega(t)=\int_{\Omega}\nabla\times\vec{u}(\cdot,t)\,\rm{d}\Omega=\int_{\partial\Omega_{N}}\!\!\!\!u_{y}(\cdot,t)\,\rm{d}s,$ (8)
@@
As noted earlier, classical eigenvalue analysis provides no information about the size of asymmetric perturbation needed to initiate a transition to an asymmetric flow profile from an unstable symmetric flow pattern. Herein, to replicate a practical laboratory experiment we will assess the symmetry or otherwise of an evolving flow generated with a distorted steady-state parabolic inlet profile,
@@
3.1. Informed data test results
[…]
@@ 3.1. Informed data test results
[…]
The data-splitting test results are unsurprising. It is clear that simpler strategies such as a $k$ nearest neighbour or a shallow decision tree classifier would be equally effective ways of visualising the decision boundary in cases like this where one is prepared to generate results with a rich set of well-chosen training points. The attractive features of the shallow neural network classifier lie in its simplicity (it is essentially parameter-free: the architecture is fixed and the optimisation strategy does not need to be specially tuned), its practicality (the raw output of the network is the probability of the bifurcation being observed), its generalisability (to high-dimensional parameter vectors), and its speed (the total time taken to complete the training was less than 4 minutes when the open-source Python code was run on a 2018 MacBook Pro).
@@
Figure 7. Symmetric step channel flow: refined grid classification (left) and prediction accuracy (right). Green dots on the right are correct predictions, blue dot is incorrect.
To assess the sensitivity to discretisation parameters and validate the surrogate model generated by the 500 data points, a second set of 50 classification data points were randomly generated close to the classification boundary with a refined spatial grid (level 6 in ifiss). In these calculations, the dimension $n_{x}$} of the algebraic system is increased from 23,842 to 94,146 and $n_{t}$}, the number of time steps, is increased from 1200 to 1600. The results are shown in Figure 9. It is reassuring to observe that only one parameter combination was misclassified by the surrogate model. The results in the next section demonstrate that data points with perturbations of magnitude between $10^{-3}$ and $10^{-4}$ prove to be the most difficult to classify correctly.
3.2. Blind data test results
[…]
Figure 8. Symmetric step channel flow showing surrogate predictions from trained neural network using 105 uniformly spaced data points. Colour variation highlights the bifurcation boundary solution (left plot). Estimated probability of observing a bifurcated solution for a fixed perturbation (right plot).
[…]